Razar Sharp Company purchased equipment on July 1, 2014, for $43,470. The equipment was expected to have a useful life of three years, or 6,480 operating hours, and a residual value of $1.350. The equipment was used for 1,200 hours during 2014, 2,300 hours in 201 s, i,900 hours in 2016, and ?,0b0 hours in 2017. Required Determine the amount of depreciation expense for the years ended December 31, 2014, 2015, 2016, and 2017, by (a) the straight-line method, (b) units-of output method and (c) the double-declining-balance method Note: FOR DECLINING BALANCE ONLY, round the multiplier to four decimal places. Then round the answer for each year to the nearest whole dollar a. Straight-line method Amount Year 2014 2015 2016 2017 Amount 2014 2015 2016 2017 Amount Year 2014
media%2F23f%2F23ffabfc-1455-497d-987a-43
Answer #1

(1) Under straight line method ........ we have Depreication per year = (Cost - salvage ) / Life in years

= (43470 - 1350) / 3 = 14040.

But during year - 1 and year - 4 we have only half year of operation, hence depreciations are as under

Year Amount
2014 7020
2015 14040
2016 14040
2017 7020

UNITS OF OUT PUT METHOD

Here we calculate the depreciation rate per unit of production ( i.e per hour )

= ( cost of asset - residual ) / Estimated life time in hours = (43470 - 1350) / 6480 hours = 6.50 per hour

Now multiply each year of hours produced with this 6.50

Year Amount
2014 7800
2015 14950
2016 12350
2017 7020

DOUBLE DECLINING METHOD

In this method, we calculate rate of Depreciation = 1 / Life * 200 = 1/3 * 200 = 66.67% or simply use 2/3 as depreciation portion. Here we should not consider the residual value.

2014 ............. 43,470 * 2/3 = 28980. But we used it for half year only = 28980 * 1/2 = 14490

2015 ........ (43470 - 14490) * 2/3 = 28980 * 2/3 = 19320

2016 .......... (28980 - 19320) * 2/3 = 9660 * 2/3 = 6440

2017 ........... (9660 - 6440) * 2/3 = 2146.67